THE FOLLOWING ARE THE HEIGHTS OF FOUR STUDENTS IN CENTIMETERS. SUPPOSE SAMPLES SIZE 2 ARE TAKEN FROM THIS POPULATION OF FOUR STUDENTS.


STUDENTS HEIGHT (in cm)


CARDO 125


ALYANA 120


JOAQUIN 130


FLORA 110



a. how many samples are possible?


b. construct the sampling distribution of the sample means


c. construct the histogram of the sampling distribuyion of the sample means

1
Expert's answer
2022-04-11T07:25:00-0400

a. "C^2_4=\frac{4!}{2!2!}=6"

b.m (125,120)=(125+120)/2=122.5

m(125,130)=(125+130)/2=127.5

m(125,110)=(125+110)/2=117.5

m(120,130)=(120+130)/2=125

m(120,110)=(120+110)/2=115

m(130,110)=(130+110)/2=120


Frequency

F(122.5)=F(127.5)=F(117.5)=F(125)=F(115)=F(120)=1

Probability

"P(x)=F(x)/\sum F(x)"

P(122.5)=P(127.5)=P(117.5)=P(125)=P(115)=P(120)=1/6

"E(x)=\sum Px=1/6(122.5+127.5+117.5+125+115+120)=121.25"

"\sigma^2=\sum Px^2-(\sum Px)^2=1/6(15006.25+16256.25+13806.25+15625+13225+14400)-14701.56=18.23"

c.